# Training Large Neural Networks With Low-Dimensional Error Feedback

### v2

@@ Abstract
Conventional wisdom holds that training large neural networks requires propagating precise, high-dimensional error gradients through every neuron, a process with limited biological plausibility. Our work demonstrates that neural networks can be trained effectively using low-dimensional error signals, comparable to the inherent complexity of the task rather than the size of the network. By developing a novel, biologically plausible learning paradigm based on adaptive low-dimensional feedback, we show theoretically and empirically that minimal error dimensionality can achieve competitive performance even in complex neural architectures. These findings challenge existing assumptions about neural optimization and suggest fundamental constraints that may shape learning in both artificial and biological systems.
@@ 1.   Introduction
Modern deep learning methods rely on high-dimensional gradients that propagate backward through layers, ensuring a precise credit assignment to each parameter [1, 2]. However, this process is computationally intensive and biologically implausible, given the lack of direct evidence for backpropagation in the brain [3, 4]. If the loss gradient does not propagate back through the network, it must be obtained through alternative routes [5]. In this case, there is no reason to assume that each neuron receives detailed information about the local gradient. Moreover, the brain likely operates with low-dimensional error signals, reflecting the inherently low-dimensional nature of tasks. This raises an essential question: Can large neural networks be trained efficiently with constrained, low-dimensional error signals? If error dimensionality is tied to task complexity rather than network size, it could revolutionize how we train large networks and improve our understanding of biological learning.
Despite the importance of this question, the effect of low-dimensional error feedback on learning and neural representations was never properly explored. The strong coupling of feedforward and backward passes in backpropagation, where both processes use the same synapses, has limited investigations into alternative training methods. While approaches such as Feedback Alignment have been proposed [5], the necessity of full gradients for effective learning has not been questioned.
@@ 3.1 Learning dynamics
Figure 2: Learning dynamics and component alignment in linear networks. (a) Schematic of the network architecture with input dimension $n=128$, hidden layer size $k=64$, and output dimension $m=64$. The feedback matrix $B$ is factorized as $B=QP$ and constrained to rank $r$. (b, c) Comparison of theoretical predictions (dashed) and numerical simulations (solid) for low-rank Feedback Alignment (FA), training only $Q$[10], with $r=64$ and $r=8$, respectively. The $y$-axis shows the alignment of singular vectors between $W_{2}W_{1}$ and $\Sigma_{io}$. (d, e) Same as (b, c) but training both and $Q$ and $P$ using Eqs (7) and (8) . (f) When $P$ is not trained, the singular modes align on average (bold), but the top $r$ components fail to recover fully. (g) Training $P$ to align with the error ensures full recovery of the top $r$ singular components, improving convergence and alignment.
@@
The implications of the uderdetrimed fixed point solutions are surprising. Naively, one would expect that if the bottleneck is not too narrow, the projections of the error would maintain the necessary structure to guide the learning. Specifically, the Johnson–Lindenstrauss lemma  [17] suggests that as long as $r>d\log m$, the pairwise correlation structure of the error signal would be maintained. Nevertheless, our analysis shows that the solutions weight are not guaranteed to converge to the correct solution. Thus, in the case of low-rank feedback, Feedback Alignment (FA) is likely to fail. The solution to this predicament, as we show next, is aligning the feedback weights with the data, ensuring that the network learns the correct representations.
@@ 5.   Error dimensionality shapes neural receptive fields
Consistent with our hypothesis, constraining the feedback rank led to the emergence of center-surround receptive fields in the retinal layer. To validate our hypothesis, we trained a model that included the feedforward bottleneck, similar to [29] but used dLDFA to train the retinal layer without restricting the backward pathway. In line with our expectations, the retinal receptive fields exhibited orientation selectivity (see the appendix).
This experiment demonstrates that error dimensionality influences neuronal tuning and neural representations. Specifically, lower-dimensional error signals promote higher symmetries in emergent receptive fields. Our findings underscore the importance of considering the dimensionality and pathways of error signals when studying neural computations in the brain.

### v3

@@ Abstract
Conventional wisdom holds that training large neural networks requires propagating precise, high-dimensional error gradients through every neuron—a process lacking biological plausibility. Our work demonstrates that effective learning can instead occur with low-dimensional error signals, whose dimensionality scales with task complexity rather than network size. By developing a novel learning paradigm that employs local plasticity rules to adaptively align feedback signals with task-relevant errors, we show theoretically and empirically that even minimal error dimensionality achieves competitive performance in complex neural architectures. These findings challenge current assumptions about neural optimization and highlight fundamental constraints that can govern learning processes in both artificial and biological neural systems.
@@ 1.   Introduction
Modern deep learning methods rely on high-dimensional gradients that propagate backward through layers, ensuring a precise credit assignment to each parameter [1, 2]. However, this process is computationally intensive and biologically implausible, given the lack of direct evidence for backpropagation in the brain [3, 4]. If the loss gradient does not propagate back through the network, it must be obtained through alternative routes [5]. In this case, there is no reason to assume that each neuron receives detailed information about the local gradient. Moreover, the brain likely operates with low-dimensional error signals, reflecting the inherently low-dimensional nature of tasks. This raises an essential question: Can large neural networks be trained efficiently with partial information about the loss gradient carried by low-dimensional error signals? If error dimensionality is tied to task complexity rather than network size, it could revolutionize how we train large networks and improve our understanding of biological learning.
Despite the importance of this question, the effect of low-dimensional error feedback on learning and neural representations was never properly explored. The strong coupling of feedforward and backward passes in backpropagation, where both processes use the same synapses, has limited investigations into alternative training methods. Some approaches, such as Feedback Alignment [5], were proposed to address the problem of weight transposition in backporpagation, the necessity of full gradients for effective learning has not been questioned.
@@ 3.1 Learning dynamics
Figure 2: Learning dynamics and component alignment in linear networks. (a) Schematic of the network architecture with input dimension $n=128$, hidden layer size $k=64$, and output dimension $m=64$. The feedback matrix $B$ is factorized as $B=QP$ and constrained to rank $r$. (b, c) Comparison of theoretical predictions (dashed) and numerical simulations (solid) for low-rank Feedback Alignment (FA), training only $Q$[10], with $r=m=64$ and $r=8<m$, respectively. The $y$-axis shows the alignment of singular vectors between $W_{2}W_{1}$ and $\Sigma_{io}$. (d, e) Same as (b, c) but training both and $Q$ and $P$ using Eqs (7) and (8) . (f) When $P$ is not trained, the singular modes align on average (bold), but the top $r$ components fail to recover fully. (g) Training $P$ to align with the error ensures full recovery of the top $r$ singular components, improving convergence and alignment.
@@
The implications of the uderdetrimed fixed point solutions are surprising. Naively, one would expect that if the bottleneck is not too narrow, the projections of the error would maintain the necessary structure to guide the learning. Specifically, the Johnson–Lindenstrauss lemma  [17] suggests that as long as $r>d\log m$, the pairwise correlation structure of the error signal would be maintained. Nevertheless, our analysis shows that the solutions weights are not guaranteed to converge to the correct solution. Thus, in the case of low-rank feedback, Feedback Alignment (FA) is likely to fail. Furthermore, FA fails when $k<m$ even if the rank of $B$ is not restricted. The solution to this predicament, as we show next, is aligning the feedback weights with the data, ensuring that the network learns the correct representations.
@@ 5.   Error dimensionality shapes neural receptive fields
Consistent with our hypothesis, constraining the feedback rank led to the emergence of center-surround receptive fields in the retinal layer. To validate our hypothesis, we trained a model that included the feedforward bottleneck, similar to [29] but used dLDFA to train the retinal layer without restricting the backward pathway. In line with our expectations, the retinal receptive fields exhibited orientation selectivity (see the appendix). This experiment demonstrates that error dimensionality influences neuronal tuning and neural representations. Specifically, lower-dimensional error signals promote higher symmetries in emergent receptive fields.
Importantly, the architecture studied by Lindsey et al. is not a faithful model of the early visual system. In particular, there is no evidence that error signals propagate from the cortex or thalamus back to the retina. Instead, it is believed that the receptive fields in the retina emerge through unsupervised learning [31]. Nevertheless, our results provide an explanation for the receptive fields observed by Lindsey et al. Furthermore, it may lead to an alternative explanation of learning in the visual pathway that uses simplified signaling (e.g., neuromodulators). Most importantly, our findings show a relation between tuning curves and error pathways.